### v1

# Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification

Abstract
Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence. This classification not only supports educational diagnostics and analytics but also enhances complex tasks like information retrieval and question answering by associating questions with relevant categories. Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance. To address this, we propose a novel approach leveraging graph convolutional networks (GCNs), named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions. By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows GCNs to learn from the interconnected nature of language more effectively. Additionally, we explore the incorporation of phrase-based features to enhance classification accuracy, especially in low-resource settings. Our findings demonstrate that GCNs, augmented with these features, offer a promising solution for more accurate and context-aware question classification, bridging the gap between graph neural network research and practical educational applications.
**footnotetext: Equal contribution
Introduction
Question classification is vital in providing AI-driven education and training, assisting adaptive learning systems by categorizing questions into skill area or level of difficulty and competence, and providing educational diagnosis and analytics. Accurate automatic question classification also provides headway to more complex tasks such as information retrieval and question answering, in which the association of the given question with a certain category allows for retrieval of relevant contextual knowledge to formulate compatible answers. Traditional methods, relying on word embeddings and conventional classifiers, often struggle with capturing the nuanced relationships between words and phrases, leading to suboptimal classification performance. The complexity of natural language, especially in the context of educational and domain-specific questions, requires more sophisticated approaches that can understand and leverage the inherent structure of language.
Aside from text embeddings, graphs have recently been highlighted as a way to represent unstructured text data. Questions can be naturally represented as graphs, where nodes correspond to words or phrases, and edges capture the relationships between them, such as syntactic dependencies, semantic similarities, or proximity measures. This graph-based representation aligns perfectly with the strengths of graph convolutional network (GCN), which excels at learning from structured data. GCNs are designed to operate on graph-structured data, making them ideal for processing and understanding the interconnected nature of words and phrases within questions. This capability enables GCNs to capture the underlying structure and dependencies in a question, leading to more accurate and context-aware classification decisions. GCNs offer a promising solution by modeling the relationships between words, phrases, and their syntactic and semantic connections as graphs. This approach allows for a more nuanced and context-aware classification, bridging the gap between basic AI research and practical applications in education and beyond.
Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts. While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification. In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
Related Work
Automatic Question Classification
One of the earliest works on Question Classification was in (Li and Roth 2002) which made use of support vector machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features. Another reference work by (Huang, Thint, and Qin 2008) also discusses usage of head words and hypernyms as features for SVM and maximum entropy (ME) algorithms. They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification (Hovy, Hermjakob, and Ravichandran 2002) via string matching in questions have also been explored in earlier works.
Following these works, a review of automatic question classifiers from 2012 to 2017 (Silva, Bittencourt, and Maldonado 2019) revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods. The approaches range from concatenating pre-trained word embeddings to form a question embedding (Godea and Nielsen 2018), to combining CNN and LSTM in a hybrid model (Zhou et al. 2016/12), effectively balancing feature extraction and sequence modeling, and was shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
Text Classification Using Text Graphs
TextGCN (Yao, Mao, and Luo 2019) pioneered approaching text classification as a node classification problem. Yao et al. constructs a corpus-level graph with document nodes and word nodes—pointwise mutual information (PMI) as word-word edge weights and text frequency-inverse document frequency (TF-IDF) as word-document edge weights—and applies a 2-layer GCN to obtain the class output. Several works propose similar graph construction and node initialization based on TextGCN, with more efficient or better regularized graph propagation (Wu et al. 2019; Zhu and Koniusz 2021).
Other works have also explored multi-graphs instead of a single corpus-level graph. TensorGCN (Liu et al. 2020) builds a tensor graph from three separate graphs: semantic graph via word embeddings, syntactic graph via grammar parsing, and sequential graph via PMI values. Each of these graphs has different word-word edge weights while sharing the same TF-IDF values for word-document edge weights, and they undergo intra- and inter-graph propagation. ME-GCN (Wang et al. 2022) constructs a graph that has multi-dimensional word-word, word-document, and document-document edges, using Word2Vec and Doc2Vec embeddings and similarity scores. SHINE (Wang et al. 2021) uses
While node classification is a transductive approach, where one needs to build a corpus-level graph consisting both train and test document nodes, InducT-GCN (Wang, Han, and Poon 2022) extends the task into a inductive approach, where a training corpus graph is constructed first similarly to TextGCN, and any new test corpus is built as a virtual subgraph for inference.
Graph classification, in which you classify a document graph as opposed to a document node, has also been used as an inductive approach to text classification. Text-Level-GNN (Huang et al. 2019) uses a sliding window to build multiple graphs, each with a small number of nodes and edges, with trainable edge weights that are shared across the graphs for the same word pair. Huang et al. also uses a message passing mechanism instead of a GCN architecture, relying on aggregating neighbourhood information via max-pooling and combining them via weighted sums, instead of convolution functions. Graph Fusion Network (Dai et al. 2022) constructs four individual global word co-occurrence graphs, and each document is represented as with a set of four subgraphs. GCN is applied to each subgraph, and the resulting output is concatenated, passed through an MLP layer, and average-pooled to obtain the final document embedding.
There are also combinatory methods such as HyperGAT (Hypergraph Attention Networks) (Ding et al. 2020) building hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling. Attention mechanism is utilized at both node- and edge-level to update representations.
Why Use Phrase Features?
As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks. We identify a research gap in adapting GCNs for question classification in the lack of question classification-specific feature extraction and graph construction methods. While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations. The effectiveness of phrases have been discussed in topic modeling for questions (Supraja, Khong, and Tatinati 2021; Supraja and Khong 2024), and this section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
Capturing Meaningful Chunks of Information
Unlike individual words, phrases (e.g., noun phrases and verb phrases) often encapsulate more specific and contextually relevant information. Consider the question ”How do enzymes function in the human body?” Here, the phrase ”enzymes function” (verb phrase) and ”human body” (noun phrase) are pivotal in identifying this as a biological process question, likely classified under ”Biology” or ”Physiology.” The verb phrase indicates the action or process in question, while the noun phrase specifies the domain of interest. In the question ”Describe the steps to solve a quadratic equation,” the phrases ”steps to solve” (verb phrase) and ”quadratic equation” (noun phrase) are linked by their relevance to problem-solving in mathematics.
This also applies to question analysis with part-of-speech (POS) tags. POS tags help in identifying the grammatical structure, but they do not capture the full semantic context. For example, knowing that ”role” is a noun does not reveal its specific relevance in a question about ”electric circuits.”
Intra-Question Relationships
By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question. For instance, in ”What are the causes of the French Revolution?”, the proximity and relationship between ”causes” and ”French Revolution” (via PPMI or other metrics) are key to categorizing this question under ”History” or ”Political Science.”
Disambiguation and Contextualization
Phrases provide additional context that can help disambiguate words that might have multiple meanings. For instance, in ”What is the role of the cell in an electric circuit?”, the noun phrase ”role of the cell” and its proximity to ”electric circuit” helps to clarify that ”cell” refers to a component of an electric circuit rather than a biological cell.
While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive and contextually rich understanding. By extracting noun phrases and verb phrases and analyzing their semantic embeddings and relationships, one can capture the nuanced meaning and intent of questions, leading to more accurate and effective classification models.
Methodology
Feature Extraction and Graph Construction
Given a corpus of question-label pair, we perform basic text cleaning which includes removal of contractions and punctuation.
To construct meaningful graph representations from a corpus of question text, we selected several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data. A summary of the extracted features are provided in Table 1.
Word PPMI Word POS Phrase PPMI Phrase POS Named Entities
Node Type Word POS tag Phrase Phrase POS tag Named entity
Node Feature Word2Vec - PhraseBERT - TransE
Edge Weight PMI value PMI value PMI value PMI value Cosine similarity
Table 1: Summary of Extracted Features
Figure 1: Process flow of feature extraction, graph construction, and model training
Words and PMI
We first tokenize the given question corpus by whitespace to obtain a unique set of words. A word-level graph is constructed, with word nodes connected by edges weighted with PMI values, representing local co-occurrence relationships. Specifically, we utilize positive PMI (PPMI), given by max(PMI($word_{1},word_{2}$), 0). Pre-trained word embeddings from word2vec (Mikolov et al. 2013) are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
Word-Level POS Tags
POS tags for each word is obtained via the default set from NLTK ***https://nltk.org and PPMI values between each pair of POS tags are calculated as edge weights for a POS tag graph.
Phrases and PMI
Phrases are extracted from each question text via regex matching for verb phrases and noun phrases with spaCy ***https://spacy.io in Python, and a unique set of phrases in the corpus is obtained. Each question is represented as a sequence of verb phrases and noun phrases present in the text. Then a phrase-level graph (similar to the word-level graph) is constructed with PMI-weighted edges. Phrase embeddings are obtained from PhraseBERT (Wang, Thompson, and Iyyer 2021) to provide additional semantic information.
Phrase-Level POS Tags
During the phrase extraction process above, spaCy also provides word-level POS tags for each word in the extracted phrases. These word-level POS tags are concatenated to form phrase-level POS tags. For example, in the question ”Discuss the main objective of layout design rules”, using the regex (ADJ)(NOUN), the noun phrase [main objective] is extracted, together with the phrase-level POS tag of (ADJ_NOUN). Then, the PPMI values between each pair of phrase-level POS tags are calculated as edge weights for a phrase-level POS tag graph.
Named Entity Recognition (NER)
Named entities (e.g., people, organizations, locations) carry significant meaning and often denote key elements within a question. A list of named entities is obtained from the NELL knowledge base (Carlson et al. 2010), and is used to extract named entities via string matching. With the short text nature of questions, it is difficult to obtain multiple entities from a single question to calculate co-occurrence statistics. Hence, we utilize TransE embeddings (Bordes et al. 2013) for each named entity and calculate the cosine similarity between embeddings of each pair of named entities, to be used as edge weights in constructing a named entity graph.
For all 5 graphs constructed above, the node features are initialized as one-hot vectors. By combining these diverse feature extraction methods, we construct a comprehensive and multi-faceted graph representation of the question text, capturing various semantic, syntactic, and named entity information crucial for question classification.
Model Metric NU ARC LREC Bloom TREC-coarse TREC-fine
CNN F1 0.085 0.226 0.198 0.752 0.782 0.627
Precision 0.048 0.171 0.141 0.763 0.783 0.653
Recall 0.333 0.333 0.333 0.750 0.790 0.644
Bi-LSTM F1 0.607 0.564 0.480 0.425 0.653 0.102
Precision 0.606 0.554 0.511 0.465 0.692 0.086
Recall 0.620 0.598 0.495 0.422 0.642 0.109
TextGCN (Yao, Mao, and Luo 2019) F1 0.722 0.694 0.671 0.663 0.730 0.505
Precision 0.715 0.687 0.677 0.652 0.675 0.500
Recall 0.735 0.735 0.679 0.680 0.780 0.589
Text-Level-GNN (Huang et al. 2019) F1 0.185 0.200 0.404 0.287 0.623 0.186
Precision 0.694 0.362 0.424 0.250 0.651 0.216
Recall 0.183 0.275 0.424 0.344 0.606 0.184
HyperGAT (Ding et al. 2020) F1 0.715 0.372 0.745 0.173 0.678 0.579
Precision 0.733 0.386 0.766 0.176 0.673 0.590
Recall 0.703 0.374 0.737 0.178 0.691 0.626
TensorGCN (Liu et al. 2020) F1 0.748 0.700 0.717 0.114 0.140 0.005
Precision 0.742 0.690 0.716 0.184 0.176 0.025
Recall 0.757 0.716 0.719 0.185 0.145 0.071
SHINE (Wang et al. 2021) F1 0.560 0.610 0.620 0.459 0.560 0.197
Precision 0.553 0.600 0.620 0.461 0.568 0.208
Recall 0.583 0.623 0.627 0.473 0.613 0.242
ME-GCN (Wang et al. 2022) F1 0.632 0.603 0.601 0.607 0.659 0.669
Precision 0.738 0.601 0.607 0.629 0.718 0.729
Recall 0.619 0.620 0.608 0.592 0.634 0.642
InducT-GCN (Wang, Han, and Poon 2022) F1 0.661 0.667 0.649 0.533 0.688 0.165
Precision 0.738 0.649 0.648 0.754 0.753 0.209
Recall 0.639 0.711 0.654 0.478 0.671 0.173
PQ-GCN F1 0.724 0.712 0.751 0.672 0.801 0.201
Precision 0.723 0.695 0.754 0.692 0.882 0.224
Recall 0.727 0.750 0.749 0.662 0.777 0.257
Table 2: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
Model Architecture
As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as be able to consolidate the each graph of different node and edge types into a single question embedding. Initial approach was to concatenate of the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function. Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
Following the initial exploration, we base our output layers largely on SHINE (Wang et al. 2021), where a dynamic graph of question nodes and cosine similarity edges is learned from the individual GCN outputs and the nodes are classified via label propagation in a final GCN layer. The training parameters are kept the same as provided by Wang et al.. An overview of the feature extraction, graph construction, and model is provided in Figure 1.
Experiments
Datasets
NU
This dataset was obtained from Najran University (Yahya et al. 2013), containing 596 questions with 3 class labels of different cognitive complexities.
ARC
This dataset was released as part of the A12 Reasoning Challenge (Boratko et al. 2018), consisting 279 questions labeled with 3 different reasoning capabilities.
LREC
This is a collection of 345 middle school science questions, labeled with 3 different type of answers required by the respective questions (Godea and Nielsen 2018).
Bloom
Bloom’s Taxonomy (Bloom and Krathwohl 1956) is one of the most well-established question classification frameworks used by educators. This dataset is a set of questions collected from different web sources, and manually classified into the six cognitive levels of Bloom’s Taxonomy (Yahya 2011).
TREC
The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC (Hovy et al. 2001), about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set. These questions were manually labeled into 6 coarse and 50 fine class labels.
Aside from TREC, for which the dataset split was obtained from HuggingFace ***https://huggingface.co/datasets/CogComp/trec, the train/test split was given by a pre-determined seed.
Performance Evaluation
We compare the performance of our proposed model with selected baseline models in Table 2, including vanilla CNN and Bi-LSTM for non-graph classification models.
PQ-GCN outperforms the baseline models in the macro-averaged evaluation on ARC, LREC, and TREC-coarse datasets, and second to the best performing model for NU and Bloom datasets. The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed by Wang et al., showing the effectiveness of phrase-based features in improving question classification capabilities.
While performance on TREC-fine drops compared to its coarse label counterpart across all models is expected due to the large number of labels, our proposed model suffers from a larger difference compared to models such as HyperGAT or ME-GCN. TREC-fine was extensively labeled and was included in experiments to determine model capability in a more extreme and granular question classification task. In realistic educational settings, it is unlikely to have a question set that is labeled as intently as TREC-fine.
Considerations for System Implementation
In this section, we outline possible considerations for practical deployment of the trained model.
Compared with Transformer-based models (Vaswani et al. 2017) which scale quadratically with sequence length, GCNs offer a less computationally extensive alternative to text embedding-based classification. This allows for greater accessibility to educators without access to appropriate compute resources.
In addition, node classification-based question classification is quick to train as at every epoch, the corpus-level graph is passed through the network once to obtain the logits for all samples, unlike traditional neural networks where samples are passed into the network batch by batch. This allows for a quick refresh of the model based on any new data points or even a whole new dataset.
There is also no need for a lot of labels for label propagation via cosine similarity at the question-level dynamic graph—some may not have access to well-labeled set of questions nor the resources for manual labeling. This is a crucial factor for learning analytics systems, and the fact that the proposed model performs well in low-resource settings shows its potential in such environments.
The proposed features can be made modular to suit the question corpus in context i.e. each individual graph can be added or removed from the model input to build the dynamic question-level graph. This will allow for greater flexibility in deployment of the model in various settings.
Conclusion and Future Works
The proposed method on applying GCNs to question classification shows promising results in capturing the intricate word and phrase relationships within question text. The next steps involve refining the model for specific educational and AI-driven applications, such as adaptive learning platforms or automated tutoring systems. Future work will focus on optimizing the model for deployment at scale, ensuring they can handle large volumes of data and provide real-time classification on user-provided questions. In conclusion, we hope to provide sufficient basis for further research into feature extraction methods for graph neural networks, especially in automatic question classification.

### v2

@@
 Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance.
-To address this, we propose a novel approach leveraging graph convolutional networks (GCNs), named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions.
-By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows GCNs to learn from the interconnected nature of language more effectively.
-Additionally, we explore the incorporation of phrase-based features to enhance classification accuracy, especially in low-resource settings.
-Our findings demonstrate that GCNs, augmented with these features, offer a promising solution for more accurate and context-aware question classification, bridging the gap between graph neural network research and practical educational applications.
-**footnotetext: Equal contribution
-Introduction
+To address this, we propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions.
+By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows the model to learn from the interconnected nature of language more effectively.
+Additionally, we explore the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics.
+Our findings demonstrate that the proposed model, augmented with these features, offer a promising solution for more robust and context-aware question classification, bridging the gap between graph neural network research and practical educational applications of AI.
+Index Terms:
+natural language processing, graph convolutional network, question classification
+11footnotetext: Equal contribution
+I Introduction
 Question classification is vital in providing AI-driven education and training, assisting adaptive learning systems by categorizing questions into skill area or level of difficulty and competence, and providing educational diagnosis and analytics.
@@
 While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification.
+Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [1] and [2] which are reviewed in Section II-A.
 In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
-Related Work
-Automatic Question Classification
-One of the earliest works on Question Classification was in (Li and Roth 2002) which made use of support vector machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features.
-Another reference work by (Huang, Thint, and Qin 2008) also discusses usage of head words and hypernyms as features for SVM and maximum entropy (ME) algorithms.
-They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification (Hovy, Hermjakob, and Ravichandran 2002) via string matching in questions have also been explored in earlier works.
-Following these works, a review of automatic question classifiers from 2012 to 2017 (Silva, Bittencourt, and Maldonado 2019) revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods.
-The approaches range from concatenating pre-trained word embeddings to form a question embedding (Godea and Nielsen 2018), to combining CNN and LSTM in a hybrid model (Zhou et al.
-2016/12), effectively balancing feature extraction and sequence modeling, and was shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
-Text Classification Using Text Graphs
-TextGCN (Yao, Mao, and Luo 2019) pioneered approaching text classification as a node classification problem.
+II Related Work
+II-A Automatic Question Classification
+One of the earliest works on Question Classification was in [3] which made use of support vector machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features.
+Another reference work by [4] also discusses usage of head words and hypernyms as features for SVM and maximum entropy (ME) algorithms.
+They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification [5] via string matching in questions have also been explored in earlier works.
+Following these works, a review of automatic question classifiers from 2012 to 2017 [6] revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods.
+The approaches include concatenating pre-trained word embeddings to form a question embedding [7], incorporating subword information in character-based convolutional neural network (CNN) [8], and combining CNN and long short-term memory (LSTM) in a hybrid model [9].
+These neural networks, effectively balancing feature extraction and sequence modeling, were shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
+While there are other works which focus on question classification using neural networks, we review the two works up to date that specifically makes use of text graphs and highlight the improvements we propose with PQ-GCN.
+The first work [1] incorporates dependency parsing between words, pointwise mutual information (PMI) between regular expressions, cosine similarity between phrase embeddings, and KL divergence between topics in a heterogeneous graph for question and short document classification.
+The authors based the architecture on TensorGCN [10] and 4 heterogeneous node sets and adjacency matrices, while PQ-GCN is based on SHINE [11] and uses 5 homogeneous node sets and adjacency matrices which reduces sparsity in input data.
+The second work [2] uses word graph based on dependency parsing, and trains a graph attention network (GAT) [12].
+This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of a question, given their short lengths.
+II-B Text Classification Using Text Graphs
+TextGCN [13] pioneered approaching text classification as a node classification problem.
 Yao et al.
-constructs a corpus-level graph with document nodes and word nodes—pointwise mutual information (PMI) as word-word edge weights and text frequency-inverse document frequency (TF-IDF) as word-document edge weights—and applies a 2-layer GCN to obtain the class output.
-Several works propose similar graph construction and node initialization based on TextGCN, with more efficient or better regularized graph propagation (Wu et al.
-2019; Zhu and Koniusz 2021).
+constructs a corpus-level graph with document nodes and word nodes—PMI as word-word edge weights and text frequency-inverse document frequency (TF-IDF) as word-document edge weights—and applies a 2-layer GCN to obtain the class output.
+Several works propose similar graph construction and node initialization based on TextGCN, with more efficient or better regularized graph propagation [14, 15].
 Other works have also explored multi-graphs instead of a single corpus-level graph.
-TensorGCN (Liu et al.
-2020) builds a tensor graph from three separate graphs: semantic graph via word embeddings, syntactic graph via grammar parsing, and sequential graph via PMI values.
+TensorGCN [10] builds a tensor graph from three separate graphs: semantic graph via word embeddings, syntactic graph via grammar parsing, and sequential graph via PMI values.
 Each of these graphs has different word-word edge weights while sharing the same TF-IDF values for word-document edge weights, and they undergo intra- and inter-graph propagation.
-ME-GCN (Wang et al.
-2022) constructs a graph that has multi-dimensional word-word, word-document, and document-document edges, using Word2Vec and Doc2Vec embeddings and similarity scores.
-SHINE (Wang et al.
-2021) uses
-While node classification is a transductive approach, where one needs to build a corpus-level graph consisting both train and test document nodes, InducT-GCN (Wang, Han, and Poon 2022) extends the task into a inductive approach, where a training corpus graph is constructed first similarly to TextGCN, and any new test corpus is built as a virtual subgraph for inference.
+ME-GCN [16] constructs a graph that has multi-dimensional word-word, word-document, and document-document edges, using Word2Vec and Doc2Vec embeddings for node features, and normalized embedding distance and TF-IDF for edge features.
+SHINE [11] uses word graphs based on three word-level relationships—PMI, POS tag co-occurrence, and cosine similarities between embeddings of entity pairs appearing in the corpus—which are easy to obtain at a low cost using common natural language processing techniques.
+While node classification is a transductive approach, where one needs to build a corpus-level graph consisting both train and test document nodes, InducT-GCN [17] extends the task into a inductive approach, where a training corpus graph is constructed first similarly to TextGCN, and any new test corpus is built as a virtual subgraph for inference.
 Graph classification, in which you classify a document graph as opposed to a document node, has also been used as an inductive approach to text classification.
-Text-Level-GNN (Huang et al.
-2019) uses a sliding window to build multiple graphs, each with a small number of nodes and edges, with trainable edge weights that are shared across the graphs for the same word pair.
+Text-Level-GNN [18] uses a sliding window to build multiple graphs, each with a small number of nodes and edges, with trainable edge weights that are shared across the graphs for the same word pair.
 Huang et al.
-also uses a message passing mechanism instead of a GCN architecture, relying on aggregating neighbourhood information via max-pooling and combining them via weighted sums, instead of convolution functions.
-Graph Fusion Network (Dai et al.
-2022) constructs four individual global word co-occurrence graphs, and each document is represented as with a set of four subgraphs.
+also uses a message passing mechanism instead of a GCN architecture, relying on aggregating neighborhood information via max-pooling and combining them via weighted sums, instead of convolution functions.
+Graph Fusion Network [19] constructs four individual global word co-occurrence graphs, and each document is represented as with a set of four subgraphs.
 GCN is applied to each subgraph, and the resulting output is concatenated, passed through an MLP layer, and average-pooled to obtain the final document embedding.
-There are also combinatory methods such as HyperGAT (Hypergraph Attention Networks) (Ding et al.
-2020) building hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling.
+There are also combinatory methods based on GAT.
+HyperGAT (Hypergraph Attention Networks) [20] builds hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling.
 Attention mechanism is utilized at both node- and edge-level to update representations.
-Why Use Phrase Features?
+Contrastive learning has also been utilized in conjunction with topic modeling and GAT to improve classification performance on short texts [21].
+III Why Use Phrase Features?
 As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks.
@@
 While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations.
-The effectiveness of phrases have been discussed in topic modeling for questions (Supraja, Khong, and Tatinati 2021; Supraja and Khong 2024), and this section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
+The effectiveness of phrases in providing deeper semantic connections in the text has been discussed in topic modeling for questions [22, 1], as well as in graph-based document filtering tasks [23].
+This section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
+Figure 1: Corpus-level graph construction with extracted features
 Capturing Meaningful Chunks of Information
-Unlike individual words, phrases (e.g., noun phrases and verb phrases) often encapsulate more specific and contextually relevant information.
-Consider the question ”How do enzymes function in the human body?” Here, the phrase ”enzymes function” (verb phrase) and ”human body” (noun phrase) are pivotal in identifying this as a biological process question, likely classified under ”Biology” or ”Physiology.” The verb phrase indicates the action or process in question, while the noun phrase specifies the domain of interest.
-In the question ”Describe the steps to solve a quadratic equation,” the phrases ”steps to solve” (verb phrase) and ”quadratic equation” (noun phrase) are linked by their relevance to problem-solving in mathematics.
+Unlike individual words, phrases (e.g.
+noun phrases and verb phrases) often encapsulate more specific and contextually relevant information.
+Consider the question “How do enzymes function in the human body?” The phrase “enzymes function” and “human body” are pivotal in identifying this as a biological process question, likely classified under “Biology” or “Physiology.” The phrases indicates the action or process in question, and also specifies the domain of interest.
+In the question “Describe the steps to solve a quadratic equation,” the phrases “steps to solve” and “quadratic equation” are linked by their relevance to problem-solving in mathematics.
 This also applies to question analysis with part-of-speech (POS) tags.
-POS tags help in identifying the grammatical structure, but they do not capture the full semantic context.
-For example, knowing that ”role” is a noun does not reveal its specific relevance in a question about ”electric circuits.”
+POS tags help in identifying the grammatical structure, but they do not capture the full semantic context when observed individually.
 Intra-Question Relationships
 By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question.
-For instance, in ”What are the causes of the French Revolution?”, the proximity and relationship between ”causes” and ”French Revolution” (via PPMI or other metrics) are key to categorizing this question under ”History” or ”Political Science.”
+For instance, in “What are the primary causes of the French Revolution?”, the proximity and relationship between “primary causes” and “French Revolution” are key to categorizing this question under “History” or “Political Science.”
 Disambiguation and Contextualization
 Phrases provide additional context that can help disambiguate words that might have multiple meanings.
-For instance, in ”What is the role of the cell in an electric circuit?”, the noun phrase ”role of the cell” and its proximity to ”electric circuit” helps to clarify that ”cell” refers to a component of an electric circuit rather than a biological cell.
+For instance, in “What is the role of the cell in an electric circuit?”, the phrase “role of the cell” and its proximity to “electric circuit” helps to clarify that “cell” refers to a component of an electric circuit rather than a biological cell.
 While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive and contextually rich understanding.
 By extracting noun phrases and verb phrases and analyzing their semantic embeddings and relationships, one can capture the nuanced meaning and intent of questions, leading to more accurate and effective classification models.
-Methodology
-Feature Extraction and Graph Construction
+IV Methodology
+IV-A Feature Extraction and Graph Construction
 Given a corpus of question-label pair, we perform basic text cleaning which includes removal of contractions and punctuation.
-To construct meaningful graph representations from a corpus of question text, we selected several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data.
-A summary of the extracted features are provided in Table 1.
-Word PPMI Word POS Phrase PPMI Phrase POS Named Entities
-Node Type Word POS tag Phrase Phrase POS tag Named entity
-Node Feature Word2Vec - PhraseBERT - TransE
-Edge Weight PMI value PMI value PMI value PMI value Cosine similarity
-Table 1: Summary of Extracted Features
-Figure 1: Process flow of feature extraction, graph construction, and model training
+To construct meaningful graph representations from a corpus of question text, we select several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data.
+A summary of the extracted features are provided in Table I.
+TABLE I: Summary of Extracted Features
+Node Type Node Feature Edge Weight
+Word PPMI Word Word2Vec PMI value
+Word POS POS tag - PMI value
+Phrase PPMI Phrase PhraseBERT PMI value
+Phrase POS Phrase POS tag - PMI value
+Named Entities Named entity TransE Cosine similarity
+Figure 2: Overview of model architecture
 Words and PMI
@@
 A word-level graph is constructed, with word nodes connected by edges weighted with PMI values, representing local co-occurrence relationships.
-Specifically, we utilize positive PMI (PPMI), given by max(PMI($word_{1},word_{2}$), 0).
-Pre-trained word embeddings from word2vec (Mikolov et al.
-2013) are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
+Specifically, we utilize positive PMI (PPMI) with a window size of 5, given by max(PMI($word_{1},word_{2}$), 0).
+PMI is given by:
+$\text{PMI}(w_{1},w_{2})=\log\frac{P(w_{1},w_{2})}{P(w_{1})P(w_{2})}$ (1)
+where $P(w_{1})=\frac{\#W(w_{1})}{\#W}$, $P(w_{2})=\frac{\#W(w_{2})}{\#W}$, $P(w_{1},w_{2})=\frac{\#W(w_{1},w_{2})}{\#W}$.
+#$\#W(w)$ is the number of sliding windows that a word $w$ in the corpus, and #$\#W(w_{1},w_{2})$ is the number of sliding windows that words $w_{1}$ and $w_{2}$ appear together in the corpus.
+#$\#W$ is the total number of sliding windows in the corpus.
+Pre-trained word embeddings from word2vec [24] are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
+TABLE II: PQ-GCN Model Parameters
+Word Embedding Size 300
+Phrase Embedding Size 768
+Entity Embedding Size 100
+Sliding Window Size for PMI 5
+Hidden Dimension Size 200
+No. of Layers in GCNs 2
+Dropout 0.7
+Threshold for Question Cosine Similarity 2.7
+Optimizer Adam
+Learning Rate 1e-3
+Weight Decay 1e-4
 Word-Level POS Tags
-POS tags for each word is obtained via the default set from NLTK ***https://nltk.org and PPMI values between each pair of POS tags are calculated as edge weights for a POS tag graph.
+POS tags for each word is obtained via the default set from NLTK  [^1] https://nltk.org and PPMI values between each pair of POS tags are calculated as edge weights for a POS tag graph.
 Phrases and PMI
-Phrases are extracted from each question text via regex matching for verb phrases and noun phrases with spaCy ***https://spacy.io in Python, and a unique set of phrases in the corpus is obtained.
+Phrases are extracted from each question text via POS tag regex matching for noun and verb phrases with spaCy  [^2] https://spacy.io in Python, and a set of unique phrases in the corpus is obtained.
+The regex patterns for noun and verb phrases are given by (2) and (3) respectively:
+[…]
+[…]
+$(\langle\text{NOUN}\rangle|\langle\text{PROPN}\rangle\langle\text{PART}\rangle?)+$ (2)
+$\mu_{V}=$ $\langle\text{AUX}\rangle*\langle\text{ADV}\rangle*\langle\text{VERB}\rangle$ (3)
+where ????, *, and $+$ represent zero or one, zero or more, and one or more occurrences of the preceding POS tag respectively.
 Each question is represented as a sequence of verb phrases and noun phrases present in the text.
 Then a phrase-level graph (similar to the word-level graph) is constructed with PMI-weighted edges.
-Phrase embeddings are obtained from PhraseBERT (Wang, Thompson, and Iyyer 2021) to provide additional semantic information.
+Phrase embeddings are obtained from PhraseBERT [25] to provide additional semantic information.
+TABLE III: Classification Label Distributions
+Dataset NU ARC LREC Bloom TREC
+Dataset Size 596 279 344 2522 5952
+Classification Cognitive complexities Reasoning capabilities Expected answer types Educational objectives Question topics
+Avg. # of Words 10.8 24.8 16.1 14.8 10.0
+Class Labels Transfer (49.7%) Linguistic matching Very short answer (41.6%) Comprehension (38.1%) Entity (22.6%)
+Knowledge (33.5%) (47.0%) Context sensitive (36.0%) Knowledge (13.6%) Description and abstract
+Application (16.8%) Basic facts (28.0%) Answers will vary (22.4%) Evaluation (12.6%) concept (21.8%)
+Hypothetical (25.0%) Application (12.5%) Human being (21.6%)
+Analysis (12.1%) Numeric value (17.0%)
+Synthesis (11.1%) Location (15.4%)
+Abbreviation (1.6%)
+TABLE IV: CNN Model Parameters
+Parameter Layer 1 Layer 2 Layer 3 Layer 4
+Type Conv2d Conv2d Conv2d Linear
+Embedding Size 300
+Sequence Length 15
+Number of Filters 100 100 100
+Kernel Size 3x3 4x4 5x5
+Stride 1x1 1x1 1x1
+Activation Function Softmax
+Dropout Rate 0.5
+TABLE V: Bi-LSTM Model Parameters
+Parameter Layer 1 Layer 2 Layer 3
+Type Bi-LSTM Bi-LSTM Linear
+Embedding Size 300
+Hidden Dimension Size 128 128
+Activation Function Softmax
 Phrase-Level POS Tags
-During the phrase extraction process above, spaCy also provides word-level POS tags for each word in the extracted phrases.
-These word-level POS tags are concatenated to form phrase-level POS tags.
-For example, in the question ”Discuss the main objective of layout design rules”, using the regex (ADJ)(NOUN), the noun phrase [main objective] is extracted, together with the phrase-level POS tag of (ADJ_NOUN).
-Then, the PPMI values between each pair of phrase-level POS tags are calculated as edge weights for a phrase-level POS tag graph.
+[…]
 Named Entity Recognition (NER)
 Named entities (e.g., people, organizations, locations) carry significant meaning and often denote key elements within a question.
-A list of named entities is obtained from the NELL knowledge base (Carlson et al.
-2010), and is used to extract named entities via string matching.
+A list of named entities is obtained from the NELL knowledge base [26], and is used to extract named entities via string matching.
 With the short text nature of questions, it is difficult to obtain multiple entities from a single question to calculate co-occurrence statistics.
-Hence, we utilize TransE embeddings (Bordes et al.
-2013) for each named entity and calculate the cosine similarity between embeddings of each pair of named entities, to be used as edge weights in constructing a named entity graph.
-For all 5 graphs constructed above, the node features are initialized as one-hot vectors.
+Hence, we utilize TransE embeddings [27] for each named entity and calculate the cosine similarity between embeddings of each pair of named entities, to be used as edge weights in constructing a named entity graph.
+From the corpus, the above features are extracted and the respective relationships are computed to obtain edge weights, which are then used to construct the graph for each feature.
+The examples of graphs constructed are shown in Figure 1.
+For the corpus-level graphs constructed above, the node features are initialized as one-hot vectors except for entity graph, which uses the TransE embeddings directly as node features.
+The mapping of which text features are present in which question is used to generate question-level embeddings during propagation.
 By combining these diverse feature extraction methods, we construct a comprehensive and multi-faceted graph representation of the question text, capturing various semantic, syntactic, and named entity information crucial for question classification.
-Model Metric NU ARC LREC Bloom TREC-coarse TREC-fine
-CNN F1 0.085 0.226 0.198 0.752 0.782 0.627
-Precision 0.048 0.171 0.141 0.763 0.783 0.653
-Recall 0.333 0.333 0.333 0.750 0.790 0.644
-Bi-LSTM F1 0.607 0.564 0.480 0.425 0.653 0.102
-Precision 0.606 0.554 0.511 0.465 0.692 0.086
-Recall 0.620 0.598 0.495 0.422 0.642 0.109
-TextGCN (Yao, Mao, and Luo 2019) F1 0.722 0.694 0.671 0.663 0.730 0.505
-Precision 0.715 0.687 0.677 0.652 0.675 0.500
-Recall 0.735 0.735 0.679 0.680 0.780 0.589
-Text-Level-GNN (Huang et al. 2019) F1 0.185 0.200 0.404 0.287 0.623 0.186
-Precision 0.694 0.362 0.424 0.250 0.651 0.216
-Recall 0.183 0.275 0.424 0.344 0.606 0.184
-HyperGAT (Ding et al. 2020) F1 0.715 0.372 0.745 0.173 0.678 0.579
-Precision 0.733 0.386 0.766 0.176 0.673 0.590
-Recall 0.703 0.374 0.737 0.178 0.691 0.626
-TensorGCN (Liu et al. 2020) F1 0.748 0.700 0.717 0.114 0.140 0.005
-Precision 0.742 0.690 0.716 0.184 0.176 0.025
-Recall 0.757 0.716 0.719 0.185 0.145 0.071
-SHINE (Wang et al. 2021) F1 0.560 0.610 0.620 0.459 0.560 0.197
-Precision 0.553 0.600 0.620 0.461 0.568 0.208
-Recall 0.583 0.623 0.627 0.473 0.613 0.242
-ME-GCN (Wang et al. 2022) F1 0.632 0.603 0.601 0.607 0.659 0.669
-Precision 0.738 0.601 0.607 0.629 0.718 0.729
-Recall 0.619 0.620 0.608 0.592 0.634 0.642
-InducT-GCN (Wang, Han, and Poon 2022) F1 0.661 0.667 0.649 0.533 0.688 0.165
-Precision 0.738 0.649 0.648 0.754 0.753 0.209
-Recall 0.639 0.711 0.654 0.478 0.671 0.173
-PQ-GCN F1 0.724 0.712 0.751 0.672 0.801 0.201
-Precision 0.723 0.695 0.754 0.692 0.882 0.224
-Recall 0.727 0.750 0.749 0.662 0.777 0.257
-Table 2: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
-Model Architecture
-As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as be able to consolidate the each graph of different node and edge types into a single question embedding.
-Initial approach was to concatenate of the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function.
+TABLE VI: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
+Best F1, precision, and recall scores are in bold and the second best scores are underlined for each dataset.
+Model Metric NU ARC LREC Bloom TREC
+CNN F1 0.085 0.226 0.198 0.752 0.782
+Precision 0.048 0.171 0.141 0.763 0.783
+Recall 0.333 0.333 0.333 0.750 0.790
+Bi-LSTM F1 0.607 0.564 0.480 0.425 0.653
+Precision 0.606 0.554 0.511 0.465 0.692
+Recall 0.620 0.598 0.495 0.422 0.642
+TextGCN [13] F1 0.722 0.694 0.671 0.663 0.730
+Precision 0.715 0.687 0.677 0.652 0.675
+Recall 0.735 0.735 0.679 0.680 0.780
+Text-Level-GNN [18] F1 0.185 0.200 0.404 0.287 0.623
+Precision 0.694 0.362 0.424 0.250 0.651
+Recall 0.183 0.275 0.424 0.344 0.606
+HyperGAT [20] F1 0.715 0.372 0.745 0.173 0.678
+Precision 0.733 0.386 0.766 0.176 0.673
+Recall 0.703 0.374 0.737 0.178 0.691
+TensorGCN [10] F1 0.412 0.499 0.566 0.107 0.805
+Precision 0.421 0.557 0.692 0.147 0.851
+Recall 0.450 0.604 0.575 0.157 0.775
+SHINE [11] F1 0.560 0.610 0.620 0.459 0.560
+Precision 0.553 0.600 0.620 0.461 0.568
+Recall 0.583 0.623 0.627 0.473 0.613
+ME-GCN [16] F1 0.632 0.603 0.601 0.607 0.659
+Precision 0.738 0.601 0.607 0.629 0.718
+Recall 0.619 0.620 0.608 0.592 0.634
+InducT-GCN [17] F1 0.661 0.667 0.649 0.533 0.688
+Precision 0.738 0.649 0.648 0.754 0.753
+Recall 0.639 0.711 0.654 0.478 0.671
+PQ-GCN F1 0.724 0.712 0.751 0.672 0.801
+Precision 0.723 0.695 0.754 0.692 0.882
+Recall 0.727 0.750 0.749 0.662 0.777
+IV-B Model Architecture
+As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding.
+Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function.
 Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
-Following the initial exploration, we base our output layers largely on SHINE (Wang et al.
-2021), where a dynamic graph of question nodes and cosine similarity edges is learned from the individual GCN outputs and the nodes are classified via label propagation in a final GCN layer.
-The training parameters are kept the same as provided by Wang et al..
-An overview of the feature extraction, graph construction, and model is provided in Figure 1.
-Experiments
-Datasets
+Following the initial exploration, we base our output layers largely on SHINE [11].
+An overview of model architecture is provided in Figure 2.
+Each graph is passed through their own 2-layer GCN.
+Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features.
+Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification.
+We provide the model and training parameters for the proposed PQ-GCN in Table II.
+V Experiments
+V-A Datasets
+We first select a number of datasets used for question classification task to evaluate our proposed model on. The details are outlined in this subsection.
 NU
-This dataset was obtained from Najran University (Yahya et al. 2013), containing 596 questions with 3 class labels of different cognitive complexities.
+This dataset was obtained from Najran University [28], containing 596 questions with 3 class labels of different cognitive complexities.
 ARC
-This dataset was released as part of the A12 Reasoning Challenge (Boratko et al. 2018), consisting 279 questions labeled with 3 different reasoning capabilities.
+This dataset was released as part of the A12 Reasoning Challenge [29], consisting 279 questions labeled with 3 different reasoning capabilities.
 LREC
-This is a collection of 345 middle school science questions, labeled with 3 different type of answers required by the respective questions (Godea and Nielsen 2018).
+This is a subset of 344 questions from the original dataset of science questions asked by teachers in real middle school classrooms, with a proposed taxonomy by expected answer types [7].
+The subset was chosen by selecting 3 mutually exclusive class labels—’Very short answer’, ’Context sensitive’, and ’Answers will vary’.
 Bloom
-Bloom’s Taxonomy (Bloom and Krathwohl 1956) is one of the most well-established question classification frameworks used by educators.
-This dataset is a set of questions collected from different web sources, and manually classified into the six cognitive levels of Bloom’s Taxonomy (Yahya 2011).
+Bloom’s Taxonomy [30] is one of the most well-established question classification frameworks used by educators, based on different educational objectives.
+This dataset is a set of questions collected from different web sources and manually classified into the six cognitive levels of Bloom’s Taxonomy [^3] The dataset can be found at dx.doi.org/10.13140/RG.2.1.4932.3123.
 TREC
-The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC (Hovy et al.
-2001), about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set.
-These questions were manually labeled into 6 coarse and 50 fine class labels.
-Aside from TREC, for which the dataset split was obtained from HuggingFace ***https://huggingface.co/datasets/CogComp/trec, the train/test split was given by a pre-determined seed.
-Performance Evaluation
-We compare the performance of our proposed model with selected baseline models in Table 2, including vanilla CNN and Bi-LSTM for non-graph classification models.
-PQ-GCN outperforms the baseline models in the macro-averaged evaluation on ARC, LREC, and TREC-coarse datasets, and second to the best performing model for NU and Bloom datasets.
-The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed by Wang et al., showing the effectiveness of phrase-based features in improving question classification capabilities.
-While performance on TREC-fine drops compared to its coarse label counterpart across all models is expected due to the large number of labels, our proposed model suffers from a larger difference compared to models such as HyperGAT or ME-GCN.
-TREC-fine was extensively labeled and was included in experiments to determine model capability in a more extreme and granular question classification task.
-In realistic educational settings, it is unlikely to have a question set that is labeled as intently as TREC-fine.
-Considerations for System Implementation
-In this section, we outline possible considerations for practical deployment of the trained model.
-Compared with Transformer-based models (Vaswani et al.
-2017) which scale quadratically with sequence length, GCNs offer a less computationally extensive alternative to text embedding-based classification.
-This allows for greater accessibility to educators without access to appropriate compute resources.
-In addition, node classification-based question classification is quick to train as at every epoch, the corpus-level graph is passed through the network once to obtain the logits for all samples, unlike traditional neural networks where samples are passed into the network batch by batch.
-This allows for a quick refresh of the model based on any new data points or even a whole new dataset.
-There is also no need for a lot of labels for label propagation via cosine similarity at the question-level dynamic graph—some may not have access to well-labeled set of questions nor the resources for manual labeling.
-This is a crucial factor for learning analytics systems, and the fact that the proposed model performs well in low-resource settings shows its potential in such environments.
-The proposed features can be made modular to suit the question corpus in context i.e.
-each individual graph can be added or removed from the model input to build the dynamic question-level graph.
-This will allow for greater flexibility in deployment of the model in various settings.
-Conclusion and Future Works
-The proposed method on applying GCNs to question classification shows promising results in capturing the intricate word and phrase relationships within question text.
-The next steps involve refining the model for specific educational and AI-driven applications, such as adaptive learning platforms or automated tutoring systems.
-Future work will focus on optimizing the model for deployment at scale, ensuring they can handle large volumes of data and provide real-time classification on user-provided questions.
-In conclusion, we hope to provide sufficient basis for further research into feature extraction methods for graph neural networks, especially in automatic question classification.
+The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC [31], about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set.
+These questions were manually labeled into 6 class labels.
+Aside from TREC, for which the dataset split is obtained from HuggingFace  [^4] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request.
+Relevant statistics for the datasets are found in Table III.
+V-B Performance Evaluation
+TABLE VII: Keyword Analysis for Five Datasets
+Frequency is calculated by % of questions in the label set that contains the keyword.
+Dataset Label Top Occurring Keywords (Frequency)
+NU Transfer would (11.8%), compare (6.4%), story (6.4%)
+Knowledge words (12.5%), describe (9.5%), state (7.0%)
+Application construct (16%), use (14%), would (12.0%)
+ARC Linguistic matching likely (22.1%), best (16.0%), water (10.7%)
+Basic facts best(16.7%), chemical (12.8%), new (10.3%)
+Hypothetical experiment (30.0%), water (27.1%), students (20.0%)
+LREC Very short answer energy (13.3%), mass (11.9%), density (11.2%)
+Context sensitive explain (18.6%), think (12.1%), lab (11.3%)
+Answers will vary energy (28.6%), explain (20.8%), would (13.0%)
+Bloom Comprehension explain (37.8%), discuss (16.6%), describe (14.3%)
+Knowledge list (22.4%), define (16.0%), five (12.8%)
+Evaluation justify (23.6%), evaluate (19.2%), answer (10.4%)
+Application calculate (14.6%), explain (13.6%), determine (11.1%)
+Analysis differentiate (17.8%), compare (15.5%), analyze (12.5%)
+Synthesis suggest (23.7%), propose (11.83%), design (10.8%)
+TREC Entity name (9.8%), fear (4.9%), first (3.9%)
+Description and abstract concept mean (5.23%), origin (4.2%), get (2.8%)
+Human being name (11.6%), first (7.38%), president (5.8%)
+Numeric value many (32.7%), year (8.2%), much (6.1%)
+Location country (13.7%), city (12.0%), state (7.2%)
+Abbreviation stand (48.4%), abbreviation (19.0%), mean (11.6%)
+We compare the performance of our proposed model against selected baseline models in Table VI, including vanilla CNN and Bi-LSTM for non-graph classification models.
+The baseline models were selected by reviewing their contributions in graph neural networks for text classification [32], each using distinct feature engineering and graph construction methods for text data, as in Section II, as compared to works which have proposed architectural improvements for computational efficiency.
+With the exception of CNN and Bi-LSTM, for which parameters were set as outlined in Table IV and Table V, the rest of the baseline models have their model and training parameters set as provided by authors in the project repositories.
+PQ-GCN achieves the best macro-averaged F1 scores across most datasets, including NU (0.724), LREC (0.751), and TREC (0.801), while consistently delivering high precision and recall.
+The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [11], showing the effectiveness of phrase-based features in improving question classification capabilities.
+CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC.
+This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category.
+An analysis of most frequently occurring words have been included in Table VII, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords.
+For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest”, “propose”, and “design” that does not appear in questions of other classes.
+While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would”, “water”, and “energy”.
+TensorGCN performs with the highest F1 score on TREC.
+TensorGCN obtains its semantic embeddings by training an LSTM model on the fly during graph construction, and this method might be effective in generating more context-accurate embeddings with larger datasets.
+While this brings about a 0.004 point gain in F1 score over PQ-GCN, it also results in a longer graph construction time with increasing dataset size due to increased training time of the LSTM model, as compared to the proposed methodology of using pre-trained embedding models.
+Bi-LSTM has a consistently average performance across the datasets, which is expected of a simple sequence model.
+TextGCN shows apt results for NU and ARC datasets where its F1 scores are second only to PQ-GCN, but there is room for improvement in leveraging text features other than PMI and TF-IDF when constructing the graph, especially when it comes to text with niche characteristics such as questions.
+Text-Level GNN shows inconsistent poor performance across the datasets, as its learnable word-word edge weights are highly dependent on relative proximity of keywords and the ability of the message passing mechanism to disambiguate polysemous words based on word proximity.
+HyperGAT also shows inconsistent performance, possibly attributed to its word-word edge features, which are extracted by segmenting a text at sentence level and treating each sentence as a “window”, as opposed to a sliding window in PMI calculation.
+Questions in the datasets usually consist of one or two sentences, and lacks usual word-sentence-document hierarchy that HyperGAT leverages on.
+ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQGCN outperforms both models in F1 scores across all datasets.
+VI Conclusion
+We develop and evaluate PQ-GCN, adapting GCNs to question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions.
+The proposed method shows promising classification performance compared to baseline models, capturing the intricate word and phrase relationships within question text.
+Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning, augmented with artificial intelligence.
+Finding suitable feature extraction methods is crucial in tackling this task, and we hope to provide sufficient basis for further research into innovative feature extraction methods for automatic text classification with graph neural networks.

### v3

@@
-Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification
+PQ-GCN: Enhancing Text Graph Question Classification with Phrase Features
 Abstract
 Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence.
-This classification not only supports educational diagnostics and analytics but also enhances complex tasks like information retrieval and question answering by associating questions with relevant categories.
-Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance.
-To address this, we propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions.
-By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows the model to learn from the interconnected nature of language more effectively.
-Additionally, we explore the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics.
-Our findings demonstrate that the proposed model, augmented with these features, offer a promising solution for more robust and context-aware question classification, bridging the gap between graph neural network research and practical educational applications of AI.
-Index Terms:
-natural language processing, graph convolutional network, question classification
-11footnotetext: Equal contribution
-I Introduction
+It not only supports educational diagnostics and analytics but also enhances complex downstream tasks like information retrieval and question answering by associating questions with relevant categories.
+Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in question statements, leading to suboptimal performance.
+We propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN).
+Through PQ-GCN, we evaluate the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics.
+The proposed method, augmented with phrase-based features, outperform baseline graph-based methods in low-resource settings, and performs competitively against language model-based methods with a fraction of their parameter size.
+Our findings offer a possible solution for more context-aware, parameter-efficient question classification, bridging the gap between graph neural network research and its educational applications.
+1 Introduction
 Question classification is vital in providing AI-driven education and training, assisting adaptive learning systems by categorizing questions into skill area or level of difficulty and competence, and providing educational diagnosis and analytics.
@@
 Aside from text embeddings, graphs have recently been highlighted as a way to represent unstructured text data.
-Questions can be naturally represented as graphs, where nodes correspond to words or phrases, and edges capture the relationships between them, such as syntactic dependencies, semantic similarities, or proximity measures.
-This graph-based representation aligns perfectly with the strengths of graph convolutional network (GCN), which excels at learning from structured data.
-GCNs are designed to operate on graph-structured data, making them ideal for processing and understanding the interconnected nature of words and phrases within questions.
+Questions can be naturally represented as graphs, where nodes correspond to different text features, and edges capture the relationships between them, such as syntactic dependencies, semantic similarities, or proximity measures.
+This graph-based representation aligns perfectly with the strengths of Graph Convolutional Networks (GCNs), which excels at learning from structured data.
+GCNs are designed to operate on graph-structured data, making them ideal for processing and understanding the interconnected nature of text features within questions.
 This capability enables GCNs to capture the underlying structure and dependencies in a question, leading to more accurate and context-aware classification decisions.
-GCNs offer a promising solution by modeling the relationships between words, phrases, and their syntactic and semantic connections as graphs.
-This approach allows for a more nuanced and context-aware classification, bridging the gap between basic AI research and practical applications in education and beyond.
+GCNs offer a promising solution by modeling the relationships between text features and their syntactic and semantic connections as graphs, which allows for a more nuanced and context-aware classification.
 Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts.
-While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification.
-Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [1] and [2] which are reviewed in Section II-A.
-In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
-II Related Work
-II-A Automatic Question Classification
+While neither tackling the task of text classification with GCNs, nor question classification via automated systems is new, our investigation aims to determine the feasibility of adapting GCNs for question classification.
+Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [23] and [20] which are reviewed in Section 2.1.
+In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings, where there is limited labeled data available.
+2 Related Work
+2.1 Automatic Question Classification
 One of the earliest works on question classification was in [17] which made use of Support Vector Machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features.
@@
 They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification [13] via string matching in questions have also been explored in earlier works.
+Figure 1: Corpus-level graph construction with extracted features. The various features extracted from each question are consolidated at the corpus level to form each feature graph.
 Following these works, a review of automatic question classifiers from 2012 to 2017 [22] revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods.
@@
 These neural networks, effectively balancing feature extraction and sequence modeling, were shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
+More recently, Pre-trained Language Models (PLMs) and Large Language Models (LLMs) which have been trained on large amounts of text data have shown their success in natural language understanding-based tasks such as question generation and classification [2]), but questions might have domain-specific terminologies or technical jargon which may not be captured fully by such generalized models.
 While there are other works which focus on question classification using neural networks, we review the two works up to date that specifically makes use of text graphs and highlight the improvements we propose with PQ-GCN.
@@
 The second work [20] uses word graph based on dependency parsing, and trains a Graph Attention Network (GAT) [25].
-This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of a question, given their short lengths.
-II-B Text Classification Using Text Graphs
+This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of questions.
+2.2 Text Classification Using Text Graphs
 TextGCN [34] pioneered approaching text classification as a node classification problem.
@@
 Attention mechanism is utilized at both node- and edge-level to update representations.
-Contrastive learning has also been utilized in conjunction with topic modeling and GAT to improve classification performance on short texts [21].
-III Why Use Phrase Features?
-As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks.
-We identify a research gap in adapting GCNs for question classification in the lack of question classification-specific feature extraction and graph construction methods.
+While these graph-based methods focus on general document-level text classification, we identify a research gap in adapting them for question classification, with the lack of question-specific feature extraction and graph construction methods, given questions’ comparatively shorter lengths and interrogative sentence structure.
+3 Why Use Phrase Features?
+As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions’ characteristics and yet be domain-agnostic to be generalizable across different classification frameworks.
+There is growing agreement in literature that our mental lexicon contains formulaic language, including idioms, phrases, and multi-word expressions [8], and there are positive benefits from understanding text through parsing by larger, more meaningful chunks rather than individual words [1].
+There is sufficient linguistic basis to explore beyond word-level representations in improving natural language understanding capabilities of automatic systems.
 While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations.
-The effectiveness of phrases in providing deeper semantic connections in the text has been discussed in topic modeling for questions [22, 1], as well as in graph-based document filtering tasks [23].
-This section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
-Figure 1: Corpus-level graph construction with extracted features
-Capturing Meaningful Chunks of Information
-Unlike individual words, phrases (e.g.
-noun phrases and verb phrases) often encapsulate more specific and contextually relevant information.
-Consider the question “How do enzymes function in the human body?” The phrase “enzymes function” and “human body” are pivotal in identifying this as a biological process question, likely classified under “Biology” or “Physiology.” The phrases indicates the action or process in question, and also specifies the domain of interest.
-In the question “Describe the steps to solve a quadratic equation,” the phrases “steps to solve” and “quadratic equation” are linked by their relevance to problem-solving in mathematics.
-This also applies to question analysis with part-of-speech (POS) tags.
-POS tags help in identifying the grammatical structure, but they do not capture the full semantic context when observed individually.
-Intra-Question Relationships
-By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question.
-For instance, in “What are the primary causes of the French Revolution?”, the proximity and relationship between “primary causes” and “French Revolution” are key to categorizing this question under “History” or “Political Science.”
-Disambiguation and Contextualization
-Phrases provide additional context that can help disambiguate words that might have multiple meanings.
-For instance, in “What is the role of the cell in an electric circuit?”, the phrase “role of the cell” and its proximity to “electric circuit” helps to clarify that “cell” refers to a component of an electric circuit rather than a biological cell.
-While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive and contextually rich understanding.
+Li and Roth [17] discuss the use of non-overlapping chunks, extracted via a trained classifier, but not explicitly defined phrases.
+Parsing by phrases and extracting phrase features have already been proven effective in some areas of computational linguistics, specifically statistical machine translation [16], neural machine translation [30], and topic modeling for questions [24, 23].
+While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive approach to understanding the question text, that can capture meaningful, disambiguated chunks of information.
 By extracting noun phrases and verb phrases and analyzing their semantic embeddings and relationships, one can capture the nuanced meaning and intent of questions, leading to more accurate and effective classification models.
-IV Methodology
-IV-A Feature Extraction and Graph Construction
+4 Methodology
+4.1 Feature Extraction and Graph Construction
 Given a corpus of question-label pair, we perform basic text cleaning which includes removal of contractions and punctuation.
 To construct meaningful graph representations from a corpus of question text, we select several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data.
-A summary of the extracted features are provided in Table I.
-TABLE I: Summary of Extracted Features
+A summary of the extracted features are provided in Table 1.
+Table 1: Summary of Extracted Features. Node features for Word POS and Phrase POS are obtained via one-hot encoding, as no viable pre-trained embeddings are available for POS tags.
 Node Type Node Feature Edge Weight
@@
 Named Entities Named entity TransE Cosine similarity
-Figure 2: Overview of model architecture
+Figure 2: Overview of model architecture.
+Each graph is passed through their own 2-layer GCN.
+Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features.
+Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification.
 Words and PMI
@@
 Pre-trained word embeddings from word2vec [19] are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
-TABLE II: PQ-GCN Model Parameters
+Table 2: PQ-GCN Model Parameters.
+Total Parameter Size varies according to number of label classes and total number of unique word-level and phrase-level POS tags (which determines embedding size via one-hot encoding), which differ in each dataset.
+We report the maximum parameter size of the model that is evaluated on TREC dataset.
 Word Embedding Size 300
@@
 Weight Decay 1e-4
+Total Parameter Size 495,406
 Word-Level POS Tags
@@
 Phrase embeddings are obtained from PhraseBERT [29] to provide additional semantic information.
-TABLE III: Classification Label Distributions
-Dataset NU ARC LREC Bloom TREC
-Dataset Size 596 279 344 2522 5952
-Classification Cognitive complexities Reasoning capabilities Expected answer types Educational objectives Question topics
-Avg. # of Words 10.8 24.8 16.1 14.8 10.0
-Class Labels Transfer (49.7%) Linguistic matching Very short answer (41.6%) Comprehension (38.1%) Entity (22.6%)
-Knowledge (33.5%) (47.0%) Context sensitive (36.0%) Knowledge (13.6%) Description and abstract
-Application (16.8%) Basic facts (28.0%) Answers will vary (22.4%) Evaluation (12.6%) concept (21.8%)
-Hypothetical (25.0%) Application (12.5%) Human being (21.6%)
-Analysis (12.1%) Numeric value (17.0%)
-Synthesis (11.1%) Location (15.4%)
-Abbreviation (1.6%)
-TABLE IV: CNN Model Parameters
-Parameter Layer 1 Layer 2 Layer 3 Layer 4
-Type Conv2d Conv2d Conv2d Linear
-Embedding Size 300
-Sequence Length 15
-Number of Filters 100 100 100
-Kernel Size 3x3 4x4 5x5
-Stride 1x1 1x1 1x1
-Activation Function Softmax
-Dropout Rate 0.5
-TABLE V: Bi-LSTM Model Parameters
-Parameter Layer 1 Layer 2 Layer 3
-Type Bi-LSTM Bi-LSTM Linear
-Embedding Size 300
-Hidden Dimension Size 128 128
-Activation Function Softmax
 Phrase-Level POS Tags
@@
 By combining these diverse feature extraction methods, we construct a comprehensive and multi-faceted graph representation of the question text, capturing various semantic, syntactic, and named entity information crucial for question classification.
-TABLE VI: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
-Best F1, precision, and recall scores are in bold and the second best scores are underlined for each dataset.
-Model Metric NU ARC LREC Bloom TREC
-CNN F1 0.085 0.226 0.198 0.752 0.782
+4.2 Model Architecture
+As we develop phrase-based features to bridge the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding.
+Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function.
+Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
+Following the initial exploration, we base our output layers largely on SHINE [31].
+An overview of model architecture is provided in Figure 2, and after a parameter sweep across different layer types, we provide the model and training parameters for the proposed PQ-GCN in Table 2.
+A training iteration is described in Algorithm 1
+Algorithm 1 Pass-through Algorithm for PQ-GCN
+1:$\{(V^{(i)},A^{(i)})\}_{i=1}^{5}$: Node sets and adjacency matrices for constructed graphs
+2:$\text{GCN}_{i}$: Two-layer GCN for the $i$-th graph
+3:$\text{GCN}_{Q}$: Two-layer GCN for question-level graph
+4:$f_{\text{linear}}$: Final linear layer for classification
+5:Predicted labels for each question
+6:for $i=1$ to $5$ do
+7:     $H^{(i)}\leftarrow\text{GCN}_{i}(V^{(i)},A^{(i)})$
+8:end for
+9:$H_{Q}\leftarrow\text{Concat}(H^{(1)},H^{(2)},\dots,H^{(5)})$
+10:$A_{Q}\leftarrow\text{CosineSimilarity}(H_{Q},H_{Q})$
+11:$Z_{Q}\leftarrow\text{GCN}_{Q}(H_{Q},A_{Q})$
+12:$\hat{Y}\leftarrow f_{\text{linear}}(Z_{Q})$
+The intermediate states $H^{(i)}$, $H_{Q}$, $A_{Q}$, and $Z_{Q}$ refer to GCN output for graph $i$, question embeddings, question-level dynamic adjacency matrix, and GCN output for question-level graph respectively.
+Table 3: Classification label distributions across evaluation datasets
+Dataset NU ARC LREC Bloom TREC
+Dataset Size 596 279 344 2522 5952
+Classification Cognitive complexities Reasoning capabilities Expected answer types Educational objectives Question topics
+Avg. # of Words 10.8 24.8 16.1 14.8 10.0
+Class Labels Transfer (49.7%) Linguistic matching Very short answer (41.6%) Comprehension (38.1%) Entity (22.6%)
+Knowledge (33.5%) (47.0%) Context sensitive (36.0%) Knowledge (13.6%) Description and abstract
+Application (16.8%) Basic facts (28.0%) Answers will vary (22.4%) Evaluation (12.6%) concept (21.8%)
+Hypothetical (25.0%) Application (12.5%) Human being (21.6%)
+Analysis (12.1%) Numeric value (17.0%)
+Synthesis (11.1%) Location (15.4%)
+Abbreviation (1.6%)
+5 Experiments
+5.1 Datasets
+We first select a number of datasets used for question classification task to evaluate our proposed model on.
+NU
+This dataset was obtained from Najran University [33], containing 596 questions with 3 class labels of different cognitive complexities.
+ARC
+This dataset was released as part of the A12 Reasoning Challenge [4], consisting 279 questions labeled with 3 different reasoning capabilities.
+LREC
+This is a subset of 344 questions from the original dataset of science questions asked by teachers in real middle school classrooms, with a proposed taxonomy by expected answer types [11].
+The subset was chosen by selecting 3 mutually exclusive class labels—’Very short answer’, ’Context sensitive’, and ’Answers will vary’.
+Bloom
+Bloom’s Taxonomy [3] is one of the most well-established question classification frameworks used by educators, based on different educational objectives.
+This dataset is a set of questions collected from different web sources and manually classified into the six cognitive levels of Bloom’s Taxonomy [^6] The dataset can be found at dx.doi.org/10.13140/RG.2.1.4932.3123.
+TREC
+The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC [12], about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set.
+These questions were manually labeled into 6 class labels.
+Aside from TREC, for which the dataset split is obtained from HuggingFace  [^7] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request.
+Relevant statistics for the datasets are found in Table 3.
+5.2 Baseline Models
+We then compare the performance of our proposed model against selected baseline models in Table 6.
+We include vanilla CNN and Bi-LSTM (with parameters in Table 4 and Table 5) for non-graph classification models; fine-tuned BERT (bert-base-uncased [^8] https://huggingface.co/google-bert/bert-base-uncased) for pre-trained language model.
+The baseline graph models were selected by reviewing their contributions in graph neural networks for text classification [28], each using distinct feature engineering and graph construction methods for text data, as in Section 2, as compared to works which have proposed architectural improvements for computational efficiency.
+With the exception of CNN, Bi-LSTM, and BERT, the rest of the baseline models have their model and training parameters set as provided by original authors in their project repositories.
+Table 4: CNN Model Parameters
+Parameter Layer 1 Layer 2 Layer 3 Layer 4
+Type Conv2d Conv2d Conv2d Linear
+Embedding Size 300
+Sequence Length 15
+Number of Filters 100 100 100
+Kernel Size 3x3 4x4 5x5
+Stride 1x1 1x1 1x1
+Activation Function Softmax
+Dropout Rate 0.5
+Table 5: Bi-LSTM Model Parameters
+Parameter Layer 1 Layer 2 Layer 3
+Type Bi-LSTM Bi-LSTM Linear
+Embedding Size 300
+Hidden Dimension Size 128 128
+Activation Function Softmax
+Table 6: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets. Best F1 scores are in bold and the second best F1 scores are underlined for each dataset.
+Type Model Metric NU ARC LREC Bloom TREC
+Non-graph models CNN F1 0.085 0.226 0.198 0.752 0.782
 Precision 0.048 0.171 0.141 0.763 0.783
@@
 Recall 0.620 0.598 0.495 0.422 0.642
-TextGCN [13] F1 0.722 0.694 0.671 0.663 0.730
+Pre-trained language model BERT F1 0.703 0.674 0.553 0.829 0.970
+Precision 0.828 0.777 0.763 0.846 0.981
+Recall 0.680 0.717 0.597 0.818 0.961
+Graph models TextGCN [34] F1 0.722 0.694 0.671 0.663 0.730
 Precision 0.715 0.687 0.677 0.652 0.675
@@
 Recall 0.727 0.750 0.749 0.662 0.777
-IV-B Model Architecture
-As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding.
-Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function.
-Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
-Following the initial exploration, we base our output layers largely on SHINE [11].
-An overview of model architecture is provided in Figure 2.
-Each graph is passed through their own 2-layer GCN.
-Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features.
-Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification.
-We provide the model and training parameters for the proposed PQ-GCN in Table II.
-V Experiments
-V-A Datasets
-We first select a number of datasets used for question classification task to evaluate our proposed model on. The details are outlined in this subsection.
-NU
-This dataset was obtained from Najran University [28], containing 596 questions with 3 class labels of different cognitive complexities.
-ARC
-This dataset was released as part of the A12 Reasoning Challenge [29], consisting 279 questions labeled with 3 different reasoning capabilities.
-LREC
-This is a subset of 344 questions from the original dataset of science questions asked by teachers in real middle school classrooms, with a proposed taxonomy by expected answer types [7].
-The subset was chosen by selecting 3 mutually exclusive class labels—’Very short answer’, ’Context sensitive’, and ’Answers will vary’.
-Bloom
-Bloom’s Taxonomy [30] is one of the most well-established question classification frameworks used by educators, based on different educational objectives.
-This dataset is a set of questions collected from different web sources and manually classified into the six cognitive levels of Bloom’s Taxonomy [^3] The dataset can be found at dx.doi.org/10.13140/RG.2.1.4932.3123.
-TREC
-The Text REtrieval Conference (TREC) dataset contains $\sim$6000 questions from four sources—4,500 English questions published by USC [31], about 500 manually constructed questions for a few rare classes, 894 TREC 8 and TREC 9 questions, and also 500 questions from TREC 10 which serves as the test set.
-These questions were manually labeled into 6 class labels.
-Aside from TREC, for which the dataset split is obtained from HuggingFace  [^4] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request.
-Relevant statistics for the datasets are found in Table III.
-V-B Performance Evaluation
-TABLE VII: Keyword Analysis for Five Datasets
-Frequency is calculated by % of questions in the label set that contains the keyword.
+We also perform an ablation study with respect to the proposed phrase and phrase POS tag features.
+For each experiment, the feature is excluded during embedding concatenation, and the rest of the model architecture remains as it is.
+The results are shown in Table 7.
+Table 7: Macro F1 scores obtained from ablation study
+Feature Removed NU ARC LREC Bloom TREC
+None (All Features) 0.724 0.712 0.751 0.672 0.801
+Phrases 0.722 0.741 0.733 0.699 0.876
+Phrase POS Tags 0.708 0.740 0.771 0.688 0.804
+Phrases & Phrase POS Tags 0.560 0.610 0.620 0.459 0.560
+6 Analysis
+PQ-GCN achieves the best macro-averaged F1 scores across NU (0.724), ARC (0.712), and LREC (0.751), while consistently delivering high precision and recall.
+The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [31], showing the effectiveness of phrase-based features in improving question classification capabilities.
+CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC.
+This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category.
+An analysis of most frequently occurring words have been included in Table 8, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords.
+For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest", “propose", and “design" that does not appear in questions of other classes.
+While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would", “water", and “energy".
+Table 8: Keyword Analysis for Five Datasets. Frequency is calculated by % of questions in the label set that contains the keyword.
 Dataset Label Top Occurring Keywords (Frequency)
@@
 Abbreviation stand (48.4%), abbreviation (19.0%), mean (11.6%)
-We compare the performance of our proposed model against selected baseline models in Table VI, including vanilla CNN and Bi-LSTM for non-graph classification models.
-The baseline models were selected by reviewing their contributions in graph neural networks for text classification [32], each using distinct feature engineering and graph construction methods for text data, as in Section II, as compared to works which have proposed architectural improvements for computational efficiency.
-With the exception of CNN and Bi-LSTM, for which parameters were set as outlined in Table IV and Table V, the rest of the baseline models have their model and training parameters set as provided by authors in the project repositories.
-PQ-GCN achieves the best macro-averaged F1 scores across most datasets, including NU (0.724), LREC (0.751), and TREC (0.801), while consistently delivering high precision and recall.
-The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [11], showing the effectiveness of phrase-based features in improving question classification capabilities.
-CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC.
-This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category.
-An analysis of most frequently occurring words have been included in Table VII, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords.
-For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest”, “propose”, and “design” that does not appear in questions of other classes.
-While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would”, “water”, and “energy”.
-TensorGCN performs with the highest F1 score on TREC.
+PQ-GCN performs well against BERT as well, especially for NU, ARC and LREC.
+However, on Bloom and TREC, BERT outperforms PQ-GCN, possibly due to the difference in the size of the dataset available for fine-tuning the model.
+TensorGCN performs with the highest F1 score on TREC among graph models.
 TensorGCN obtains its semantic embeddings by training an LSTM model on the fly during graph construction, and this method might be effective in generating more context-accurate embeddings with larger datasets.
@@
 Questions in the datasets usually consist of one or two sentences, and lacks usual word-sentence-document hierarchy that HyperGAT leverages on.
-ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQGCN outperforms both models in F1 scores across all datasets.
-VI Conclusion
-We develop and evaluate PQ-GCN, adapting GCNs to question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions.
-The proposed method shows promising classification performance compared to baseline models, capturing the intricate word and phrase relationships within question text.
-Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning, augmented with artificial intelligence.
-Finding suitable feature extraction methods is crucial in tackling this task, and we hope to provide sufficient basis for further research into innovative feature extraction methods for automatic text classification with graph neural networks.
+ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQ-GCN outperforms both models in F1 scores across all datasets.
+The results of the ablation study conducted in Table 7 shows that adding either phrase-based feature graph into training the model improves upon the base model, while their combined effectiveness varies across datasets.
+This variation is expected as the proposed approach extracts noun and verb phrases dynamically for each dataset using regex patterns, and the resulting phrase set and hence constructed graph differs for each dataset.
+7 Limitations
+PQ-GCN adopts a transductive approach to learning the representations, and hence it is not suited for on-line inference.
+While the focus of this work is to improve graph-based question classification with phrase features, we hope to inspire future work in incorporating phrase-based features into inductive methods.
+We evaluate PQ-GCN only on English-based question datasets.
+For the proposed method to be applied to other languages, the feature extraction methods must first be adapted to syntactic rules of the target language.
+However, extracting regex-based noun or verb phrases may be challenging for non-configurational languages with free word order [7].
+We also evaluate PQ-GCN against BERT—a representative encoder-based language model—for a more complete comparison, but not decoder-based modern large language models, as fine-tuning these models is prohibitively expensive, and generative models are not aligned with our larger focus on evaluating additive text features in embedding-based classification.
+We highlight that our method is more effective in low-resource settings compared to high-resource settings.
+However, labeling of a large question bank is resource-intensive for educators, and they are more likely to have a smaller subset of questions to be analyzed, in which case our proposed method would be succinct.
+8 Conclusion
+We develop and evaluate PQ-GCN, which augments a SHINE-based GCN for question classification by incorporating phrase features into the text graphs that can provide deeper semantic connections within questions.
+The proposed method shows competitive classification performance compared to baseline models, especially in low-resource settings without the need for extensive pre-training.
+Question classification is an important task in building educational analytics systems, as we move towards an era of personalized learning with artificial intelligence.
+Finding suitable feature extraction methods is crucial in tackling downstream tasks in an increasingly data-scarce landscape, and our proposed method provides a parameter- and data-efficient approach to enhancing question classification.